Find the marginal profit function if cost and revenue are given by C(x) = 255 +0.8x and R(x) = 7x -0.02x2. P(x)=0
Answer #1

solution : Given that ; Cost CCD) - 255 +0.8 x Revenue R(x) = 78-0.022² Profit RS) - CC1) = (x - 6.02 x Tx -0.02x²-255 _0.88